OpenStudy (anonymous):

What does counting rate mean in the below question? please help.. A radioactive source, having a half life of 5 minutes, was found to have a counting rate of 12000 counts/sec ata certain point in time. What is the counting rate fifteen minutes after this observation?

OpenStudy (anonymous):

Counting rate implies rate of nuclear disintegration. Some elements are unstable and they disintegrate into daughter nuclei accompanied with the emission of radiation (photons). For each photon hit to the the respective detector surface a "HIT" is registered and this adds to the count. 12000 counts/sec means the source nuclei sample is disintegrating into 12000 daughter nuclei per second. Hope you got it.

OpenStudy (anonymous):

after 5 minutes, half of the original material has decayed into something that is presumably stable.

OpenStudy (anonymous):

to be noted... radioactive disintegration can be accompanied with emission of particles like alpha, beta, gamma (photons) or others, depending upon the process.

OpenStudy (anonymous):

Thanx a lot!

OpenStudy (anonymous):

here...i found this online..its a a little different then ur problem..but i hope this will help..it gives u a step-by-step thing : http://answers.yahoo.com/question/index?qid=20090317095329AA1RdK2

OpenStudy (anonymous):

After 5 minutes it would be half the count 12000/2 = 6000 after ten minutes it would be half that again, 6000/2 = 3000, after 15 minutes it would be half that again, 3000/2 = 1500 counts. (or {2 x 2 x 2 = 8} 12000/8 = 1500 counts)